perimeter of a rectangle is 260 cm. the width is 15cm less than the length, what are the dimensions of the rectangle?
Answer by checkley77(12844)

P=2L+2W
W=L-15
260=2L+2(L-15)
260=2L+2L-30
4L=260+30
4L=290
L=290/4
L=72.5 ANS. FOR THE LENGTH.
W=72.5-15=57.5 ANS. FOR THE WIDTH.
PROOF:
260=2*72.5+2*57.5
260=145+115
260=260